OpenStudy (anonymous):

Simplify. Write answer with positive exponents (15x)(x^3)/(24x^-2)

OpenStudy (campbell_st):

5/8 x^6

OpenStudy (anonymous):

Thats not an option on my hw

OpenStudy (campbell_st):

lol... well the exponent is x^ (1 + 3 - (-2))= x^6 15/24 simplifies to 5/8... put them together

OpenStudy (anonymous):

Could it be 5x^6/8 ?

OpenStudy (campbell_st):

lol.. yes its \[5/8 \times x^6\]